OpenStudy (anonymous):

2one-fourth r = 10 what does r = in this equation?

OpenStudy (texaschic101):

2 1/4 r = 10 lets turn our mixed number into an improper fraction 9/4r = 10 -- divide both sides by 9/4 (9/4r) / (9/4) = 10 / (9/4) r = 10 / (9/4) r = 10 * 4/9 r = 40/9 and if you need it in mixed number form, it is 4 4/9

OpenStudy (the_fizicx99):

That is exactly what I answered last night wtf ;~;

OpenStudy (anonymous):

lolz